A one peso coin is dropped from the leaning tower of Pisa. It atarts from rest and falls freely. Compute its position and velocity after 2.0,3.0and 4.0 seconds.
1
Expert's answer
2019-01-30T15:50:08-0500

The position of a coin with respect to a tower top

s(t)=(gt^2)/2

We get

s(2s)=〖9.8×2.0〗^2/2=19.6 m

s(3s)=〖9.8×3.0〗^2/2=44.1 m

s(4s)=〖9.8×4.0〗^2/2=78.4 m

The velocity of a coin

v(t)=gt

So

v(2s)=9.8×2.0=19.6 m/s

v(3s)=9.8×3.0=29.4 m/s

v(4s)=9.8×4.0=39.2 m/s